OpenStudy (anonymous):

just started college a week ago. need some refreshing in slopes and how to put an equation in graph form. like how do you put 2x+y=4 on a graph.

OpenStudy (anonymous):

y=-2x+4

OpenStudy (anonymous):

ok, this is precisely wat i am doing right now for being a freshman in hs first, find the x and y intercepts 2(0)+y=4 + y int. 2x+ (0)=4 + x int then graph intercepts x axis is horizontal, y axis is vertical, fyi

OpenStudy (anonymous):

oops, the second + signs in those 2 equations should be = signs, my bad!